OpenStudy (anonymous):

Solve |5x + 2| = 13.

OpenStudy (jamesj):

ivan, we don't hand out answers with no explanation.

OpenStudy (anonymous):

{3, -2.2} {13, -13} {2.2, -2.2} {2.2, -3} these are the choices??

OpenStudy (anonymous):

JamesJ, dont be mad you cant solve it

OpenStudy (jamesj):

Take a step back, what is the definition of |x|?

OpenStudy (jamesj):

@ivan514, don't be absurd. Of course I can solve it.

OpenStudy (konradzuse):

ok so lets not attack the mod and lets explain. As James has said what is the definition of |x| +/-x right? |5x + 2| = 13 5x + 2 = 13 5x+2 = -13

OpenStudy (jamesj):

right. Konrad, take it from here, thanks.

OpenStudy (konradzuse):

-2 from both sides 5x = -15 5x = 11 divide both sides by 5 -15/5 = x 11/5 = x -3 = x 11/5 = x